OpenStudy (anonymous):

88 coins are worth $14.80 the coins are nickels and quarters. how many nickels and how many quarters?

OpenStudy (bobobox):

okay lets see 14.80 divided by 5 (nickels) is 269.... so we need to group nickels into groups of five until there are only 88 coins

OpenStudy (bobobox):

... right mr.billj5

OpenStudy (bobobox):

I don't know fancy math vocabulary

OpenStudy (bobobox):

oh you don't need the actual answer?

OpenStudy (bobobox):

there is only one solution

OpenStudy (bobobox):

I think

OpenStudy (anonymous):

ok, yes i don't care so much for the answer just how you set up the system?

OpenStudy (bobobox):

oh...yeah can help you there im sorry

OpenStudy (anonymous):

ah, ty, yes im very bad at algebra

OpenStudy (anonymous):

let no. of nickels=x no. of quarters=y x+y=88 ...(1) .05 x+0.25 y=14.80 5x+25y=1480 x+5y=296 ...(2) (2)-(1) gives 4y=296-88=208 y=52 x=88-52=36

OpenStudy (anonymous):

@surjithayer ty, but i dont understand this

OpenStudy (anonymous):

@surjithayer can you please explain it to me

OpenStudy (anonymous):

nvm, got it ty

OpenStudy (anonymous):

one nickel=5 cents x nickels=5x cents=0.05 x dollars one quarter=25 cents y quarters=25 y cents=0.25 y dollars now x+y=88 and 0.05 x+0.25 y=14.80

OpenStudy (anonymous):

thank you :)